### v1

# LLM-SQL-Solver: Can LLMs Determine SQL Equivalence?

Abstract
Judging the equivalence between two SQL queries is a fundamental problem with many practical applications in data management and SQL generation (i.e, evaluating the quality of generated SQL queries in text-to-SQL task). While the research community has reasoned about SQL equivalence for decades, it poses considerable difficulties and no complete solutions exist. Recently, Large Language Models (LLMs) have shown strong reasoning capability in conversation, question answering and solving mathematics challenges. In this paper, we study if LLMs can be used to determine the equivalence between SQL queries under two notions of SQL equivalence (semantic equivalence and relaxed equivalence). To assist LLMs in generating high quality responses, we present two prompting techniques: Miniature & Mull and Explain & Compare. The former technique is used to evaluate the semantic equivalence in which it asks LLMs to execute a query on a simple database instance and then explore if a counterexample exists by modifying the database. The latter technique is used to evaluate the relaxed equivalence in which it asks LLMs to explain the queries and then compare if they contain significant logical differences. Our experiments demonstrate using our techniques, LLMs is a promising tool to help data engineers in writing semantically equivalent SQL queries, however challenges still persist, and is a better metric for evaluating SQL generation than the popular execution accuracy.
1 Introduction
Structured Query Language (SQL) it is the standard programming language to query data systems and it is one of the most popular programming language  [^1] https://www.stackscale.com/blog/most-popular-programming-languages/. Since SQL is not a procedural language, one may write different SQL queries that are semantically equivalent. As a result, determining the equivalence of SQL queries is a fundamental task in many practical important applications ranging from query optimizers in database systems Pirahesh et al. (1992); Bai et al. (2023); Zhou et al. (2021) to evaluating the translation from user text-to-SQL Zelle and Mooney (1996); Zhong et al. (2020b); Khatry et al. (2023). However, it has been shown by the data management research community that checking the equivalence of two SQL queries is in general undecidable Abiteboul et al. (1995). While there are theoretical works to check equivalence under different assumptions Chandra and Merlin (1977); Sagiv and Yannakakis (1980); Cohen et al. (1999); Chu et al. (2018) and implemented prototypes such as Cosette Chu et al. (2017), they only support limited SQL operations and such projects are no longer maintained due to lack of experts. For instance, Cosette does not support commonly found operations like EXISTS, IN, queries with constant expressions, etc.; and Cosette is no longer maintained due to the lack of staff. Query optimizers and query processing engines in data systems still primarily rely on heuristic query rewriting rules Begoli et al. (2018). Exploring and writing these rules correctly require strong domain expertise and need many engineering hours.
Moreover, due to the difficulty in evaluating the equivalence of two SQL queries, Text-to-SQL literature Zhong et al. (2017); Yu et al. (2018); Lee et al. (2021); Li et al. (2023) primarily use exact match (exact SQL string match) or execution match (compare the execution outputs) as the evaluation metrics Zhong et al. (2020a). However, exact string match may overlook many equivalent SQL queries (there are many ways different ways to write the same SQL query) and execution match may misclassify different SQL queries as equivalent due to no counterexamples are contained in the underlining database instance.
Recent advancements in large language models (LLMs) have shown great capability in following instructions and in generating decisions that align with human judgement Bubeck et al. (2023); Chiang and Lee (2023); Yan et al. (2023a). For instance, It has been shown that GPT models are capable of performing first order logic reasoning  Saparov and He (2022), detect implicit contents Huang et al. (2023); Zhang et al. (2023b), and refining the invisible logic Feng et al. (2023). In this paper, we investigate whether LLM can act as a judge to determine the equivalence of SQL queries. Hence, in this paper, we would like to mainly answer the following questions:
  • Can LLMs help data engineers in identifying SQL semantic equivalence?
  • Should LLMs be used to evaluate the quality of generated SQL queries with the expected SQL answer?
2 Background
2.1 Motivation
To demonstrate the capability of LLMs, we first conduct an initial study on a real-world query rewrite error which is known as the COUNT bug Ganski and Wong (1987). Here are two SQL queries Q1 and Q2 as shown below:
1Q1:
2SELECT Parts.pnum FROM Parts
3WHERE Parts.qoh = (SELECT COUNT(Supply.shipdate)
4  FROM Supply
5  WHERE Supply.pnum = Parts.pnum
6  AND Supply.shipdate < 10);
7
8Q2:
9WITH Temp AS
10  SELECT Supply.pnum, COUNT(Supply.shipdate) AS ct
11  FROM Supply
12  WHERE Supply.shipdate < 10
13  GROUP BY Supply.pnum
14SELECT Parts.pnum FROM Parts,Temp
15WHERE Parts.qoh = Temp.ct
16AND Parts.pnum = Temp.pnum;
At first sight, one may think these two queries (the COUNT bug) are equivalent, since they both return results based on the condition that qoh from Supply table equals to the counts of records in the Supply table when shipdate less than 10; however they are different in how they handle the subquery. Suppose that there is a record $R$ in the Parts table such that R.qoh is zero and R.pnum does not exist in the Supply table. Then, Q1 will include record $R$ since R.qoh is zero and selecting the count for R.pnum from the Supply table will also return zero. On the other hand, Q2 will not include record $R$ because the TEMP result will not contain R.pnum.
We gave this equivalence question to ChatGPT and ChatGPT is able to correctly identify that these two SQL queries are not equivalent. When asked to provide a counterexample, it is able to generate a counterexample in which it provided a database instance containing such a record $R$ as described above.
While proving the semantic equivalency between two SQL queries has been proved to be undecidable Trahtenbrot (1963); Libkin (2004); Sagiv and Yannakakis (1980), we believe LLMs may offer new opportunities to help data engineers in scrutinizing their SQL optimization logic and potentially identify flaws accordingly.
2.1.1 Two Notions of SQL Equivalence
First, we define two different notions on SQL equivalence motivated from query optimizers in data systems and evaluation from the text-to-SQL literature.
Definition 2.1 (Semantic Equivalence).
Two SQL queries Q1 and Q2 are said to be semantically equivalent if and only if there does not exist a database instance such that Q1 and Q2 return different outputs.
Definition 2.2 (Relaxed Equivalence).
Two SQL queries Q1 and Q2 are said to be relaxed equivalent if one may edit a trivial portion of the SQL queries while preserving the same logic to make Q1 and Q2 semantically equivalent.
Based on the semantic definition of SQL equivalence (Definition 2.1), the Count bug is indeed a flaw for database’ query engines and should be remedied immediately. On the other hand, we argue that based on Definition 2.2, which reflects the logical perspective and aligned with human perception, the two SQL queries in the Count Bug example is relaxed logically equivalent. The fact that even experienced experts failed to identify this subtle difference indicates humans often recognize equivalence based on the relaxed notion. As a result, we believe that the relaxed notion of SQL equivalence can better capture a human user’s satisfaction with the generated SQL from the foundational models. We asked ChatGPT if the two queries in the Count Bug are relaxed equivalent and it successfully identified them as relaxed equivalent because they share the same logic.
To illustrate why these two queries are relaxed logically equivalent, Q2 in the Count Bug example only the TEMP table needs to be edited as shown below for it to be semantically equivalent to Q1 in the Count Bug example:
1Q2 TEMP Table Original:
2WITH Temp AS
3  SELECT Supply.pnum, COUNT(Supply.shipdate) AS ct
4  FROM Supply
5  WHERE Supply.shipdate < 10
6  GROUP BY Supply.pnum
7Q2 TEMP Table After Edits:
8WITH Temp AS
9  SELECT Parts.pnum, COUNT(Supply.shipdate) AS ct
10  FROM Parts, Supply
11  WHERE Supply.shipdate < 10
12  AND Parts.pnum = Supply.pnum
Figure 1: Demonstration of Counter Example, Miniature & MULL, and Explain % Compare prompts. The used SQL queries are not semantically equivalent but are relaxed equivalent.
For text-to-SQL task, we believe that when comparing the gold SQL query (expected answer) with the generated SQL query, the current most popular metric, execution match, does not reflects the human perspective and fails to capture a user’s opinion on the generated SQL. For instance, a common failure reasons in execution match is that the generated SQL selects fewer/extra columns than the gold SQL. In fact, a recent work has explicitly added "do not select extra columns" in its prompt to improve the execution accuracy Dong et al. (2023). However, from user’s perspective providing extra columns may not reduce the quality of the answer. Also, there are many other similar reasons that can lead to false execution matches but not necessarily indicate a lower quality of the generated SQL query (such as the column order in selection, case sensitivity, and duplicate entries or extra aggregates in the output). Furthermore, execution matching fails to correctly identify semantically different queries when the underlying database does not contain any counter examples, even when developer spend hundreds of hours on creating the benchmark Yu et al. (2018). As a result, it would be desirable to have a SQL solver which can determine SQL equivalence without constructing sophisticated databases and without accessing any stored data. Hence, in additional to investigating the capability of foundational models in identifying semantic equivalence, we further explore whether a foundational model can be used as a better metric for text-to-SQL generation benchmarks.
3 Proposed Method
Motivated by recent methods Deng et al. (2023); Yao et al. (2023); Wei et al. (2022), we present our prompting techniques to obtain high quality judgments from Open-AI GPT models. Our prompting technique adheres to key aspects widely adopted in NL2SQL works Li et al. (2023); Khatry et al. (2023); Liu et al. (2023); Taori et al. (2023b): 1) Clear Layout: We provide the instruction at the beginning, followed by the two SQL queries that need to be compared, and end by providing the database schema. 2) Clear Context: We provide the database schema and explicitly included the primary and foreign key constraints, since they are crucial context information for deciding SQL equivalence.
For semantic equivalence, we also added a short definition of SQL semantic equivalence (i.e., Note: Two SQL queries are semantically equivalent if and only if they yield identical results for all possible databases.). In addition, we first ask GPT to generate a counter example if there exists one and then decide if two SQL queries are equivalent. We call this prompt the Counter Example prompt. However, as shown in Figure 1, the counter example prompt leads to sub-optimal results, especially when GPT models only scratched the surface and observed the obvious (similarity between two SQL queries in structure and logic). This finding aligns with the prior works Wang et al. (2023); Zheng et al. (2023) such that GPT models exhibit bias toward ignoring small differences when determining SQL equivalence and exhibit limit reasoning capability.
To overcome the observed phenomena, we introduce our proposed prompting technique in detail to obtain better responses from large language models. In complex problem solving, people always start with a simple example and then introduce more complex elements to understand the full picture. Motivated by this simple intuition, we incorporate this idea and proposed the Miniature and Mull prompting technique. In particular, we first ask GPT to execute both queries on a simple database and then see if it can modify the database such that the two queries output different results (i.e., a counter example). As shown later in the experimental section, this approach can directly enhance the LLM’s response accuracy in determining SQL semantic equivalence. We believe that this prompting technique has wide applications in comparing the equivalence of two codes, plans, and logic.
For relaxed equivalence, we designed the Explain and Compare prompt by asking the LLMs to perform multi-step reasoning before determining if the two SQL queries are equivalent or not. This prompt follows the main ideas from the chain of thoughts Wei et al. (2022). We first prompt the LLMs to explain two SQL queries and then ask it to determine if these two queries contain significant difference in logic, as shown in Figure 1. We observe that LLMs are inclined to label two SQL queries are equivalent when they appear similar in structure. As a result, it is crucial for LLMs to be able to first spot the differences between these two queries and then apply this knowledge in considering if these two SQL queries are relaxed equivalent or not.
4 Evaluations
We evaluate the capability of popular large language models in identifying semantic and relaxed equivalence between a pair of SQL queries. In particular, we showcase the experimental results using the popular GPT 3.5 turbo (gpt-3.5-turbo) and GPT 4 turbo (gpt-4-1106-preview) models. The models are accessed via OpenAI API.
Hyper-parameter: We use zero demonstrations in our prompts. In addition, in the API call, we have set temperature to 0 which implies greedy sampling is used to generate output; The max tokens are set to 3000; and the stopping tokens are set to "Answer:" for all experiments.
4.1 Data Source
Our evaluation was conducted on i) Non-equivalent Queries Chu et al. (2017) and ii) Spider cross-domain benchmark Yu et al. (2018). The Nonequivalent Queries are originally used to evaluate the performance of Cosette a formal language based SQL solver; however, Cosette does not support many SQL operations such as IN, EXCEPT, etc.; and the Cosette project is no longer maintained. We use Nonequivalent Queries to investigate if LLMs can successfully notice the semantic differences (often subtle) between two non-equivalent SQL queries.
  • Nonequivalent Queries Chu et al. (2017): It contains 4 challenging and 1 easy questions, including the COUNT BUG, one exam question from CSE344 (data management) at the University of Washington, two bugs users found on SQL solver, and a simple hello world example.
The Spider benchmark contains 1034 development questions covering 20 databases in different application domains, and each database contains multiple tables. The Spider SQL queries are classified into four levels of complexity: easy, medium, hard, and extra. This classification is based on various factors, including the quantity of SQL keywords utilized, the incorporation of nested subqueries, and the usage of column selections and aggregations.
The official evaluation metrics on the Spider benchmark are execution accuracy and exact string match accuracy. Since two semantically equivalent SQL queries can be written in many different ways, execution accuracy is often consider as more important than exact string match. In all experiments, we use the open sourced execution match evaluation code from GitHub  [^2] https://github.com/THU-BPM/ChatGPT-sql. However, as we argued in previous sections, execution match neither captured the semantic equivalence notion nor the relaxed equivalence notion. As a result, we showcase the results of using LLMs to determine if the generated SQL queries are the equivalent or not. We used three open-sourced results.
  • Chat Liu et al. (2023): Generate SQL queries using ChatGPT with zero-shot basic prompt.
  • DIN Pourreza and Rafiei (2023): Generate SQL queries by decomposing the task into fine grained sub-modules and use the GPT-4 model with few-shot prompting
  • DAIL Gao et al. (2023): The current SOTA open sourced technique on the Spider benchmark. It uses the GPT-4 model with few-shot prompting and incorporates self-consistency Wang et al. (2022).
4.2 Finding Counter Examples
Judge Prompt Accuracy
GPT 3.5 turbo CounterExample 4/5
Miniature&Mull 5/5 $\uparrow$
GPT 4 turbo CounterExample 3/5
Miniature&Mull 5/5 $\uparrow$
GPT 4 CounterExample 4/5
Miniature&Mull 4/5 -
Table 1: Results in Nonequivalent Queries
We test LLMs ability in deciding if two SQL queries are semantically equivalent. Recall, two queries can not be semantically equivalent if there exists a counter example (a database instance) such that two queries produce different outputs.
We used two different prompting methods Counter Example and Miniature & Mull to compare and contrast the accuracy of LLMs. An answer is accurate if the answer states the two SQL queries are not equivalent.
Judge Data Source Equivalence/Prompt Easy Medium Hard Extra Acc
- Spider - 248 446 174 166 -
Execution Match Chat - 221 331 101 72 70.1%
DIN - 228 382 132 101 81.5%
DAIL - 226 392 132 104 82.6%
GPT 3.5 turbo Chat Semantic/Miniature&Mull 210 293 84 48 61.4%
DIN Semantic/Miniature&Mull 207 307 96 76 66.3%
DAIL Semantic/Miniature&Mull 224 317 85 95 70.4%
[…]
DIN Relaxed/Explain&Compare 223 360 133 114 80.3%
DAIL Relaxed/Explain&Compare 218 371 140 119 82.0%
GPT 4 turbo Chat Semantic/ Miniature&Mull 199 244 78 55 55.7%
DIN Semantic/Miniature&Mull 209 308 99 73 66.6%
DAIL Semantic/Miniature&Mull 216 343 114 75 73.0%
[…]
DIN Relaxed/Explain&Compare 209 377 124 106 79.0%
DAIL Relaxed/Explain&Compare 220 382 141 110 82.5%
Table 2: A breakdown of the accuracy scores when using different metrics, i.e., execution accuracy and LLMs judgements with varying prompt techniques. The compared SQL queries are the gold SQL queries from Spider and the generated SQL queries from Chat, DIN, and DAIL.
As shown in Table 1, when using the Counter Example prompt (asking LLMs to determine SQL equivalence and provide a counter example if two SQL queries is not equivalent), GPT 4 and GPT 4 turbo failed to notice the subtle difference in the COUNT BUG, whereas GPT 3.5 turbo correctly output that queries in COUNT BUG are not semantically equivalent. In addition, GPT 4 turbo failed on a user found bug originally posted on Hackers News  [^3] https://github.com/uwdb/Cosette/issues/29. This bug contains two SQL queries where one used DISTINCT in selection and the other used UNION ALL to union two sub-queries. However, GPT 4 turbo failed to recognize that UNION ALL will not remove duplicate rows (though UNION will). GPT 3.5 turbo failed on the CSE344 exam question which involves different sub-queries.
When using Miniature & Mull (asking LLMs to first run a simple example over the given SQL queries and then examining if the example database can be modified to be a counter example), GPT 3.5 turbo and GPT 4 turbo successfully identified all the mistakes. However, GPT 4 still failed on the COUNT BUG. We can observe that, when using Counter Example prompt (follows the main ideas of chain of thoughts), LLMs scores at least 60% accuracy. When using the Miniature & Mull prompting, the accuracy is significantly improved to 100% for GPT 3.5 turbo and GPT 4 turbo. As a result, we believe that letting LLMs execute the SQL queries on a simple database instance and then considering a counter example is an effective way to determine SQL semantic equivalence. Moreover, since GPT 4 turbo showcases strong performance with faster response time and less monetary costs [^4] https://openai.com/pricing compared to GPT 4, we use GPT 3.5 turbo and GPT 4 turbo for all the following experiments.
4.3 Judging NL2SQL Generation
There are two notions that one can take in judging the equivalence of two SQL queries, i.e., the semantic and relaxed notions. We use the proposed Miniature & Mull and Explain & Compare prompts to evaluate the equivalence of SQL queries following semantic and relaxed notions respectively.
In Table 2, we present the breakdown of the alignment between human preference and the execution match/LLM judgement on the SQL pairs (gold SQL from Spider and generated SQL from Chat, DIN, and DAIL). We observe that the relative ranking between Chat, DIN, and DAIL are preserved for both execution matches and LLM judgements. Since the generated SQL queries in Chat are from ChatGPT with zero-shot prompting, we expect the SQL queries from DIN and DAIL to achieve better accuracy. Since DAIL applied the self-consistency techniques (resource-intensive Gao et al. (2023)), we expect DAIL to be more likely to produce the best answers. As shown in Table 2, the ranking among Chat, DIN, and DAIL based on execution accuracy, GPT 3.5 turbo judgments, and GPT 4 turbo judgements with varying prompts, all aligned with our expectation.
4.3.1 Agreement between LLMs Judgements and Execution Match
We observe that for a majority of the questions, LLMs judgements and execution matches agree with each other. In particular, as shown in, Table 3, agreements between LLMs and execution match for the semantic equivalence notion have at least 69.6%; similarly, the agreements between LLMs and execution match for the relaxed equivalence notion have at least 82.4%. We believe this is an encouraging signal that implies the good quality of execution matches and the good quality of LLMs judgements. Assuming the LLMs semantic equivalence judgement is highly accurate, a high percentage of agreements indicate the execution match correctly identified many semantically different SQL queries and the underlying database contains high quality data points (up to 1000 human hours were spent on creating the Spider Benchmark Yu et al. (2018)). If the underlying database does not contain enough high quality data points, then many of the non-semantically equivalent SQL queries will return the same results. Assuming the underlying database is well designed and the execution match is highly accurate, then a high agreement between execution match and LLMs are obviously desirable.
Judge Equiv Data Agreement
GPT 3.5 turbo Semantic Chat 69.6%
DIN 72.6%
DAIL 74.6%
[…]
DIN 78.9%
DAIL 84.6%
GPT 3.5 turbo Relaxed Chat 83.8%
DIN 88.4%
DAIL 90.0%
[…]
DIN 82.4%
DAIL 86.1%
Table 3: We compare the agreement between LLMs judgements and execution matches.
4.3.2 Human Evaluation on Samples
Although, the agreements between execution accuracy and LLM’s judgments are high, there is a still lots of questions with disagreements. In this section, we provide a detailed analysis on the sampled questions from the set of questions where execution match and LLM’s judgement differs. Since there are hundreds of such questions (different results between execution match and LLM) per data source and per GPT models, we sample 10 questions randomly from the SQL query pairs where LLMs and execution match disagree.
Equivalence Evaluated Manually To evaluate the equivalence between two SQL queries, we use the human preference approach. We manually check the equivalence of the sampled questions where LLMs and execution match disagree. In human evaluation, we have also differentiated between the semantic equivalence and relaxed equivalence (in essence, semantic equivalence is objective and relaxed equivalence is subjective),
Quality Control To ensure the quality of human judgement, we have three independent SQL experts judging the equivalence and we use the majority vote to determine the label for each pair of SQL queries.
Judge Equiv Data Prefer
GPT 3.5 turbo Semantic Chat 2/10 $\downarrow$
DIN 4/10 $\downarrow$
DAIL 6/10 $\uparrow$
[…]
DIN 7/10 $\uparrow$
DAIL 5/10 -
GPT 3.5 turbo Relaxed Chat 6/10 $\uparrow$
DIN 6/10 $\uparrow$
DAIL 6/10 $\uparrow$
[…]
DIN 7/10 $\uparrow$
DAIL 9/10 $\uparrow$
Table 4: We compare the human preference between LLM judge and execution match on 10 randomly sampled questions from different data sources. The questions are selected from a set where LLMs and execution match have different opinions on these questions
Semantic Equivalence
When using the semantic equivalence notion, we find GPT 3.5 turbo may not deliver better judgment than execution match, but GPT 4 turbo showcase better alignments. When the execution match detect discrepancy in the SQL queries’ outputs, it often indicates these two SQL queries are not semantically equivalent. There are a few rare cases that we observed which the SQL queries are semantically equivalent but execution match determines not equivalent. For example, when the provided database in Spider contains column values that can not be decoded (similar issues are also raised  [^5] https://github.com/taoyds/spider/issues/46), semantically equivalent queries can be misclassified by execution match.
Figure 2: The correct operation rates for LLMs. Correct operations imply the LLM correctly executed and reasoned about the SQL queries on its own provided simple database.
In Table 4, GPT 3.5 turbo aligned with human experts on 2 out of 10 questions and 4 out of 10 questions for Chat and DIN respectively. Diving deep into the questions which GPT 3.5 turbo wrongly classified the SQL queries, we find that there are common reasons for the wrong classifications: Distinct entries vs duplicated entries. GPT 3.5 turbo often failed to consider the difference between selecting unique rows and selecting rows that may be duplicated; Subtle difference in column names. When the gold SQLs and generated SQLs share very similar SQL structure but contains subtle differences in the columns used such as group by id v.s. group by student_id, GPT 3.5 turbo may fail to notice the small difference; Hallucination. GPT 3.5 turbo may returns wrong results from executing the SQL queries on its own created simple database instance and sometimes when the outputs are indeed identical, it may still state that the outputs are different. This hallucination snowballing effect Zhang et al. (2023a) is one of the top reason why GPT 3.5 turbo perform worse compared to GPT 4 turbo.
When GPT 4 turbo is used as the judge, GPT 4 turbo constantly produced more accurate or comparable results to execution match, with at least 5 out of 10 questions for DAIL and at most 10 out of 10 questions for Chat. One interesting observation is that GPT 4 turbo correctly identified the difference between order by asc/desc limit 1 v.s. where col = select min/max(col). The former will always return one row, as limit 1 is used. The latter, however, may return many rows when these rows share the same min or max value. Execution match often failed to capture this difference. Although GPT 4 turbo judgment is much better than GPT 3.5 turbo, GPT 4 turbo may still miss on the small differences such as similar but different column names (song_name v.s. name), overlook the fact that extra columns are selected, and hallucinate. As shown in Figure 2, we looked into how well does LLMs executed and reasoned about the SQL queries on its own generated simple database. We find that GPT 4 turbo has much higher correct operation rates and hence much better response.
One common phenomena that lead to false answers in execution match is due to the return of the empty tuple. When the underlying database does not contain tuples that match the predicate, gold SQLs and generated SQLs will always match, irrespective of how different they are. Another interesting observation raised from a few comparisons is about the case sensitivity. OpenAI GPT models think databases are by default case insensitive; Spider benchmark, on the other hand, uses the SQLite databases which are case sensitive by default. Many popular database are case insensitive such as SQL Server Mistry and Misner (2014) and MySQL DuBois (2013). We believe that treating case sensitivity in string comparison as an excessive step, and therefore, we operate under the assumption that the underlying database instance is case insensitive when evaluating semantic equivalence.
Relaxed Equivalence When using the explain and compare prompt for determining the relaxed equivalence, both GPT 3.5 turbo and GPT 4 turbo consistently better aligned with human judgment compared with execution match, as shown in Table 4. GPT 3.5 turbo aligned with human preference on at least 6 out of 10 questions and GPT 4 turbo aligned with human preference on at least 7 out of 10 questions. Recall, the relaxed notion mimics the opinion of the users of NL2SQL services. The positive results on the alignment between LLMs judgments with human preferences indicate optimizing for LLMs judgments accuracy is a better optimization target than execution match accuracy.
The main reasons for why LLMs judgements are better aligned with human preference than execution match is that LLMs can zoom into the logic and determine whether the two SQL queries use the same exact logic and contain the same truth. For instance, order by asc/desc limit 1 v.s. where col = select min/max(col) may return different or the same outputs depending on the underlying database instance. Since execution match strictly checks the outputs, it may sometimes return equivalent and other times return not equivalent. However, based on the relaxed equivalence definition, order by asc/desc limit 1 is the same as where col = select min/max(col) since they both are selecting the rows with minimal or maximal value.
One interesting observation we found is about null values. GPT often consider COUNT(*), COUNT(primary key), and COUNT(DISTINCT primary key) to be different. The reason is that GPT models think that null values can be inserted into the column with primary key constraint. In fact, the SQLite databases [^6] https://www.sqlite.org/lang_createtable.html, by default, allow null values to be inserted to primarily key. However, in many database primary key constraints explicitly states that null values are not allowed [^7] https://docs.oracle.com/javadb/10.10.1.2/ref/rrefsqlj13590.html. Based on the relaxed equivalence definition, we consider records in the primary columns to be not null. Although LLMs correctly argued that when null values are presented in the primary key column, then the given SQL queries lead to different results (such as COUNT(*) considers null values as unique entries, where as COUNT(primary key) ignores rows with null values), we consider these not as a significant issue.
5 Discussion and Limitation
There are some limitations or areas of improvement to this work. First, we only evaluated the capability of Large Language Models of OpenAI. There are more other models such as Llama2 7B Touvron et al. (2023), Vicuna Chiang et al. (2023), Alpaca Taori et al. (2023a) which also showcase strong understanding and reasoning for logical tasks Chen et al. (2021); Yue et al. (2023); Yan et al. (2023b). Nevertheless, we believe that GPT 3.5 turbo and GPT 4 turbo exemplify the current status of GPT model family’s capabilities. In addition, in the manually evaluating the equivalence under two different views (semantic and relaxed), it is done over the sampled questions. Since there are more than two thousand of SQL queries per data source and equivalence view, manually checking all of them are arguably infeasible. In fact, even checking a pair of SQL queries for semantic view takes around 5 minutes for experts in SQL. Although the questions are sampled, we observe a consistent comparisons and trends between LLMs’ judgements and execution matches (such as GPT 4 turbo’s responses are consistently better aligned with human preferences).
6 Conclusion
In this paper, we studies and answer two main questions: i) Can current LLMs be used to help data engineers in checking semantic equivalences between SQL queries and ii) should we use LLMs to judge the quality of generated SQL queries for text-to-SQL benchmarks. Through extensive evaluations and manually examining the quality of the answers based on the alignment with human preferences, we find current LLMs (especially GPT 4 turbo), using our proposed prompt techniques, can assists engineers in determining semantic equivalent queries, but may yield incorrect responses. In addition, when comparing the generated SQL query with the gold SQL query, LLMs’ responses achieve high alignments with human preferences. Moreover using LLMs to determine relaxed SQL equivalence avoid the construction of sophisticated databases and no data stored need to be accessed. To be best of our knowledge, this is the first work studying the ability of LLMs in determining SQL equivalence. As the learning ability in LLMs continuously to advance, we believe that using these models to check SQL equivalence is a promising direction.

### v2

@@
 (2023); Zhou et al.
+(2021); Samwel et al.
 (2018) to evaluating the translation from user text-to-SQL Zelle and Mooney (1996); Zhong et al.
 (2020b); Khatry et al.
+(2023); Floratou et al.
 (2024).
@@
 (2017); Yu et al.
+(2018); Wang et al.
 (2019); Lee et al.
@@
 (2023); Chiang and Lee (2023); Yan et al.
-(2023a).
+(2023a); Zhang et al.
+(2023b).
 For instance, It has been shown that GPT models are capable of performing first order logic reasoning  Saparov and He (2022), detect implicit contents Huang et al.
 (2023); Zhang et al.
-(2023b), and refining the invisible logic Feng et al.
+(2023c), and refining the invisible logic Feng et al.
 (2023).
@@
 Equivalence Evaluated Manually To evaluate the equivalence between two SQL queries, we use the human preference approach.
-We manually check the equivalence of the sampled questions where LLMs and execution match disagree.
-In human evaluation, we have also differentiated between the semantic equivalence and relaxed equivalence (in essence, semantic equivalence is objective and relaxed equivalence is subjective),
-Quality Control To ensure the quality of human judgement, we have three independent SQL experts judging the equivalence and we use the majority vote to determine the label for each pair of SQL queries.
+Three of the authors manually check the equivalence of the sampled questions where LLMs and execution match disagree.
+In addition, we have also differentiated the equivalence between the semantic notion and relaxed notion (in essence, the former is objective and the latter is subjective).
+To ensure the quality of human judgement, the annotations are conducted independently, and we use the majority vote to determine the label for each pair of SQL queries (i.e., each pair of SQL queries are examined by three times independently).
 Judge Equiv Data Prefer
@@
 DIN 4/10 $\downarrow$
+DAIL 5/10 -
+[…]
+DIN 6/10 $\uparrow$
+DAIL 5/10 -
+GPT 3.5 turbo Relaxed Chat 6/10 $\uparrow$
+DIN 5/10 $\uparrow$
 DAIL 6/10 $\uparrow$
 […]
-DIN 7/10 $\uparrow$
-DAIL 5/10 -
-GPT 3.5 turbo Relaxed Chat 6/10 $\uparrow$
 DIN 6/10 $\uparrow$
-DAIL 6/10 $\uparrow$
-[…]
-DIN 7/10 $\uparrow$
-DAIL 9/10 $\uparrow$
+DAIL 7/10 $\uparrow$
 Table 4: We compare the human preference between LLM judge and execution match on 10 randomly sampled questions from different data sources.
@@
 Figure 2: The correct operation rates for LLMs. Correct operations imply the LLM correctly executed and reasoned about the SQL queries on its own provided simple database.
-In Table 4, GPT 3.5 turbo aligned with human experts on 2 out of 10 questions and 4 out of 10 questions for Chat and DIN respectively.
+In Table 4, GPT 3.5 turbo aligned with human experts in the range of two to five out of ten.
 Diving deep into the questions which GPT 3.5 turbo wrongly classified the SQL queries, we find that there are common reasons for the wrong classifications: Distinct entries vs duplicated entries.
@@
 Execution match often failed to capture this difference.
-Although GPT 4 turbo judgment is much better than GPT 3.5 turbo, GPT 4 turbo may still miss on the small differences such as similar but different column names (song_name v.s.
-name), overlook the fact that extra columns are selected, and hallucinate.
+Although GPT 4 turbo judgment is much better than GPT 3.5 turbo, GPT 4 turbo may still miss on the small differences such as similar but different column names (such as song_name v.s.
+name), overlook the fact that extra columns are selected, and also sometimes hallucinate.
 As shown in Figure 2, we looked into how well does LLMs executed and reasoned about the SQL queries on its own generated simple database.
-We find that GPT 4 turbo has much higher correct operation rates and hence much better response.
+We find that GPT 4 turbo has much higher correct operation rates which leads much better response comapred with the weaker model.
 One common phenomena that lead to false answers in execution match is due to the return of the empty tuple.
@@
 We believe that treating case sensitivity in string comparison as an excessive step, and therefore, we operate under the assumption that the underlying database instance is case insensitive when evaluating semantic equivalence.
-Relaxed Equivalence When using the explain and compare prompt for determining the relaxed equivalence, both GPT 3.5 turbo and GPT 4 turbo consistently better aligned with human judgment compared with execution match, as shown in Table 4.
-GPT 3.5 turbo aligned with human preference on at least 6 out of 10 questions and GPT 4 turbo aligned with human preference on at least 7 out of 10 questions.
-Recall, the relaxed notion mimics the opinion of the users of NL2SQL services.
+Relaxed Equivalence When using the explain and compare prompt for determining the relaxed equivalence, both GPT 3.5 turbo and GPT 4 turbo are consistently better aligned with human judgment or at par compared to execution match, as shown in Table 4.
+Recall, the relaxed notion aims to illustrate the degree to which humans perceive two queries as being similar in their meaning, and mimics the opinion of the users of NL2SQL services.
 The positive results on the alignment between LLMs judgments with human preferences indicate optimizing for LLMs judgments accuracy is a better optimization target than execution match accuracy.
-The main reasons for why LLMs judgements are better aligned with human preference than execution match is that LLMs can zoom into the logic and determine whether the two SQL queries use the same exact logic and contain the same truth.
-For instance, order by asc/desc limit 1 v.s.
-where col = select min/max(col) may return different or the same outputs depending on the underlying database instance.
+The main reasons for why LLMs judgements are better aligned with human preference than execution match is that LLMs can zoom into the query logic and determine whether the two SQL queries share the same logic and contain the same truth.
+Consider the following example, order by col asc/desc limit 1 v.s.
+where col = select min/max(col).
+They may return different or the same outputs after execution (depends on the underlying database instance).
 Since execution match strictly checks the outputs, it may sometimes return equivalent and other times return not equivalent.
-However, based on the relaxed equivalence definition, order by asc/desc limit 1 is the same as where col = select min/max(col) since they both are selecting the rows with minimal or maximal value.
+However, LLMs can correctly identify that these two queries are selecting the rows with minimal or maximal value and these two queries are equivalent under the relaxed notion,
 One interesting observation we found is about null values.
@@
 Since there are more than two thousand of SQL queries per data source and equivalence view, manually checking all of them are arguably infeasible.
-In fact, even checking a pair of SQL queries for semantic view takes around 5 minutes for experts in SQL.
-Although the questions are sampled, we observe a consistent comparisons and trends between LLMs’ judgements and execution matches (such as GPT 4 turbo’s responses are consistently better aligned with human preferences).
+In fact, even checking a pair of SQL queries for semantic view takes around 10 minutes for experts in SQL.
+Although the questions are sampled, we observe a consistent trends between LLMs’ judgements and execution matches (such as GPT 4 turbo’s responses are consistently better aligned with human preferences).
 6 Conclusion
@@
 Through extensive evaluations and manually examining the quality of the answers based on the alignment with human preferences, we find current LLMs (especially GPT 4 turbo), using our proposed prompt techniques, can assists engineers in determining semantic equivalent queries, but may yield incorrect responses.
-In addition, when comparing the generated SQL query with the gold SQL query, LLMs’ responses achieve high alignments with human preferences.
-Moreover using LLMs to determine relaxed SQL equivalence avoid the construction of sophisticated databases and no data stored need to be accessed.
-To be best of our knowledge, this is the first work studying the ability of LLMs in determining SQL equivalence.
+On the other hand, when comparing the generated SQL query with the gold SQL query with relaxed notion of equivalence, LLMs’ responses constantly achieve high alignments with human preferences.
+Moreover, another advantage of using LLMs to determine SQL equivalence is that it can avoid the construction of sophisticated databases, and no data stored need to be accessed.
+To the best of our knowledge, this is the first work studying the ability of LLMs in determining SQL equivalence.
 As the learning ability in LLMs continuously to advance, we believe that using these models to check SQL equivalence is a promising direction.

### v3

@@
 Abstract
-Judging the equivalence between two SQL queries is a fundamental problem with many practical applications in data management and SQL generation (i.e, evaluating the quality of generated SQL queries in text-to-SQL task).
+Judging the equivalence between two SQL queries is a fundamental problem with many practical applications in data management and text-to-SQL generation.
 While the research community has reasoned about SQL equivalence for decades, it poses considerable difficulties and no complete solutions exist.
-Recently, Large Language Models (LLMs) have shown strong reasoning capability in conversation, question answering and solving mathematics challenges.
+Recently, Large Language Models (LLMs) have shown strong reasoning capability in conversation, question answering and mathematics challenges.
 In this paper, we study if LLMs can be used to determine the equivalence between SQL queries under two notions of SQL equivalence (semantic equivalence and relaxed equivalence).
@@
 The latter technique is used to evaluate the relaxed equivalence in which it asks LLMs to explain the queries and then compare if they contain significant logical differences.
-Our experiments demonstrate using our techniques, LLMs is a promising tool to help data engineers in writing semantically equivalent SQL queries, however challenges still persist, and is a better metric for evaluating SQL generation than the popular execution accuracy.
+Our experiments demonstrate that using our techniques, LLMs are promising tools to help data engineers in writing semantically equivalent SQL queries, however challenges still persist, and have better alignment with human preference on text-to-SQL generations than execution accuracy.
 1 Introduction
-Structured Query Language (SQL) it is the standard programming language to query data systems and it is one of the most popular programming language  [^1] https://www.stackscale.com/blog/most-popular-programming-languages/.
+Structured Query Language (SQL) it is the standard programming language to query data systems and it is one of the most popular programming language Stackscale .
 Since SQL is not a procedural language, one may write different SQL queries that are semantically equivalent.
@@
 (2023); Floratou et al.
-(2024).
+(2024a).
 However, it has been shown by the data management research community that checking the equivalence of two SQL queries is in general undecidable Abiteboul et al.
@@
 (2021); Li et al.
-(2023) primarily use exact match (exact SQL string match) or execution match (compare the execution outputs) as the evaluation metrics Zhong et al.
+(2023) primarily use exact match (exact SQL string match) or execution match (compare the execution outputs) as the evaluation metrics.
+However, exact string match may overlook many equivalent SQL queries (there are many ways different ways to write the same SQL query), and execution match may misclassify different SQL queries as equivalent due to the lack of counterexamples in the underlining database instance.
+We note that there have been a few attempts that explore using semantic equivalence as the evaluation metrics Zhong et al.
 (2020a).
-However, exact string match may overlook many equivalent SQL queries (there are many ways different ways to write the same SQL query) and execution match may misclassify different SQL queries as equivalent due to no counterexamples are contained in the underlining database instance.
+These methods assess query semantic equivalence by generating numerous fuzzy database instances, leading to significant computational and storage demands.
+Specifically, evaluating a thousand pairs of SQL queries requires at least 75 minutes and consumes about 3 GB of storage space for creating these fuzz database instances Zhong et al.
+(2020a).
 Recent advancements in large language models (LLMs) have shown great capability in following instructions and in generating decisions that align with human judgement Bubeck et al.
@@
 (2023b).
-For instance, It has been shown that GPT models are capable of performing first order logic reasoning  Saparov and He (2022), detect implicit contents Huang et al.
+For instance, It has been shown that GPT models are capable of performing first order logic reasoning  Saparov and He (2022), detecting implicit contents Huang et al.
 (2023); Zhang et al.
-(2023c), and refining the invisible logic Feng et al.
+(2023c), and solving complex logical problems Feng et al.
+(2023); Yao et al.
 (2023).
@@
   • Can LLMs help data engineers in identifying SQL semantic equivalence?
-  • Should LLMs be used to evaluate the quality of generated SQL queries with the expected SQL answer?
+  • Should LLMs be used to evaluate text-to-SQL generation?
 2 Background
-2.1 Motivation
-To demonstrate the capability of LLMs, we first conduct an initial study on a real-world query rewrite error which is known as the COUNT bug Ganski and Wong (1987).
-Here are two SQL queries Q1 and Q2 as shown below:
+2.1 Motivations
+We first introduce a query rewriting error that happened in the real-world system, known as the COUNT bug Ganski and Wong (1987). Here are two SQL queries Q1 and Q2 as shown below:
 1Q1:
@@
 16AND Parts.pnum = Temp.pnum;
-At first sight, one may think these two queries (the COUNT bug) are equivalent, since they both return results based on the condition that qoh from Supply table equals to the counts of records in the Supply table when shipdate less than 10; however they are different in how they handle the subquery.
+One may think these two queries are equivalent, since they both return results based on the condition that qoh from Supply table equals to the counts of records in the Supply table when shipdate is less than 10; however they are different in how they handle the subquery.
 Suppose that there is a record $R$ in the Parts table such that R.qoh is zero and R.pnum does not exist in the Supply table.
@@
 On the other hand, Q2 will not include record $R$ because the TEMP result will not contain R.pnum.
-We gave this equivalence question to ChatGPT and ChatGPT is able to correctly identify that these two SQL queries are not equivalent.
-When asked to provide a counterexample, it is able to generate a counterexample in which it provided a database instance containing such a record $R$ as described above.
+Hence, real-world data systems need tools to avoid such pitfalls.
 While proving the semantic equivalency between two SQL queries has been proved to be undecidable Trahtenbrot (1963); Libkin (2004); Sagiv and Yannakakis (1980), we believe LLMs may offer new opportunities to help data engineers in scrutinizing their SQL optimization logic and potentially identify flaws accordingly.
+For instance, when we present the COUNT BUG to ChatGPT, ChatGPT is able to correctly identify that these two SQL queries are not equivalent.
+Moreover, in text-to-SQL generation, many natural language questions are inherently ambiguous Katsogiannis-Meimarakis and Koutrika (2023); Floratou et al.
+(2024b), and as a result, they can be interpreted and answered in multiple ways Min et al.
+(2020).
+From the user’s perspective, different SQL queries that convey the same logical answer to a question may all be considered equivalent.
+However, the widely used execution match metric, which expects a single correct SQL query, does not capture these ambiguities.
+Consequently, we are interested in exploring whether LLMs can be utilized to assess the quality of text-to-SQL generation.
 2.1.1 Two Notions of SQL Equivalence
-First, we define two different notions on SQL equivalence motivated from query optimizers in data systems and evaluation from the text-to-SQL literature.
+Motivated from the query optimizers in data systems and evaluation metrics for text-to-SQL, we define two notions of SQL equivalence.
 Definition 2.1 (Semantic Equivalence).
@@
 Two SQL queries Q1 and Q2 are said to be relaxed equivalent if one may edit a trivial portion of the SQL queries while preserving the same logic to make Q1 and Q2 semantically equivalent.
-Based on the semantic definition of SQL equivalence (Definition 2.1), the Count bug is indeed a flaw for database’ query engines and should be remedied immediately.
-On the other hand, we argue that based on Definition 2.2, which reflects the logical perspective and aligned with human perception, the two SQL queries in the Count Bug example is relaxed logically equivalent.
-The fact that even experienced experts failed to identify this subtle difference indicates humans often recognize equivalence based on the relaxed notion.
-As a result, we believe that the relaxed notion of SQL equivalence can better capture a human user’s satisfaction with the generated SQL from the foundational models.
-We asked ChatGPT if the two queries in the Count Bug are relaxed equivalent and it successfully identified them as relaxed equivalent because they share the same logic.
-To illustrate why these two queries are relaxed logically equivalent, Q2 in the Count Bug example only the TEMP table needs to be edited as shown below for it to be semantically equivalent to Q1 in the Count Bug example:
-1Q2 TEMP Table Original:
-2WITH Temp AS
-3  SELECT Supply.pnum, COUNT(Supply.shipdate) AS ct
-4  FROM Supply
-5  WHERE Supply.shipdate < 10
-6  GROUP BY Supply.pnum
-7Q2 TEMP Table After Edits:
-8WITH Temp AS
-9  SELECT Parts.pnum, COUNT(Supply.shipdate) AS ct
-10  FROM Parts, Supply
-11  WHERE Supply.shipdate < 10
-12  AND Parts.pnum = Supply.pnum
 Figure 1: Demonstration of Counter Example, Miniature & MULL, and Explain % Compare prompts. The used SQL queries are not semantically equivalent but are relaxed equivalent.
-For text-to-SQL task, we believe that when comparing the gold SQL query (expected answer) with the generated SQL query, the current most popular metric, execution match, does not reflects the human perspective and fails to capture a user’s opinion on the generated SQL.
-For instance, a common failure reasons in execution match is that the generated SQL selects fewer/extra columns than the gold SQL.
+2.2 Why LLM-SQL-Solver
+First, LLM-SQL-Solver (semantic notion) can be directly used to find system bugs (e.g., COUNT BUG) and help engineers validate their SQL rewriting code bases.
+Second, since query optimizer relies on heuristic rules to find equivalent SQL forms, it may not explore all rewriting possibilities and hence, the rewritten query may lead to sub optimal performance.
+We believe that, with LLM-SQL-Solver, query optimizers can explore the vast universe of equivalent rewritten queries and then find the query plan with near optimal performance.
+For text-to-SQL tasks, we believe that when comparing the gold SQL query (the correct answer) with the generated SQL query, the current most popular metric, execution match, does not reflects the human perspective and fails to capture a user’s opinion on the generated SQL.
+For instance, a common reasoning for failures in execution match is that the generated SQL selects fewer/extra columns than the gold SQL.
 In fact, a recent work has explicitly added "do not select extra columns" in its prompt to improve the execution accuracy Dong et al.
 (2023).
-However, from user’s perspective providing extra columns may not reduce the quality of the answer.
-Also, there are many other similar reasons that can lead to false execution matches but not necessarily indicate a lower quality of the generated SQL query (such as the column order in selection, case sensitivity, and duplicate entries or extra aggregates in the output).
-Furthermore, execution matching fails to correctly identify semantically different queries when the underlying database does not contain any counter examples, even when developer spend hundreds of hours on creating the benchmark Yu et al.
+However, from a user’s perspective providing extra columns may not reduce the quality of the answer.
+Also, there are many other similar reasons that can lead to false execution matches but not necessarily indicate a lower quality of the generated SQL query (such as the column order in selection, case sensitivity, duplicate entries, and columns in different tables share the similar contents).
+Moreover, execution matching fails to correctly identify logically different queries when the underlying database does not contain any counter examples, even when hundreds of human hours were consumed to create the benchmark Yu et al.
 (2018).
-As a result, it would be desirable to have a SQL solver which can determine SQL equivalence without constructing sophisticated databases and without accessing any stored data.
-Hence, in additional to investigating the capability of foundational models in identifying semantic equivalence, we further explore whether a foundational model can be used as a better metric for text-to-SQL generation benchmarks.
+Lastly, a fundamental limitation of execution accuracy is its need to query the underlying database.
+This is particularly pertinent as the text-to-SQL service provider usually does not have ownership of the user’s database and the user’s database may also contain sensitive information.
+As a result, it would be desirable to have a SQL solver determining the relax notion of equivalence without constructing sophisticated databases and without accessing any stored data.
+Hence, in additional to investigating the capability of foundational models in identifying semantic equivalence, we further explore whether LLM-SQL-Solver (relaxed notion) can be used as a better metric for text-to-SQL benchmarks.
 3 Proposed Method
@@
 (2023); Zheng et al.
-(2023) such that GPT models exhibit bias toward ignoring small differences when determining SQL equivalence and exhibit limit reasoning capability.
-To overcome the observed phenomena, we introduce our proposed prompting technique in detail to obtain better responses from large language models.
+(2023) which demonstrated that GPT models exhibit bias toward ignoring small differences when determining SQL equivalence and exhibit limited reasoning capability.
+To overcome this observed phenomenon, we introduce our proposed prompting technique in detail to obtain better responses from large language models.
 In complex problem solving, people always start with a simple example and then introduce more complex elements to understand the full picture.
@@
 We believe that this prompting technique has wide applications in comparing the equivalence of two codes, plans, and logic.
-For relaxed equivalence, we designed the Explain and Compare prompt by asking the LLMs to perform multi-step reasoning before determining if the two SQL queries are equivalent or not.
-This prompt follows the main ideas from the chain of thoughts Wei et al.
+For relaxed equivalence, we designed the Explain and Compare prompt by asking the LLMs to perform multi-step reasoning before determining if the two SQL queries are equivalent.
+This prompt follows the main ideas from the chain-of-thoughts Wei et al.
 (2022).
 We first prompt the LLMs to explain two SQL queries and then ask it to determine if these two queries contain significant difference in logic, as shown in Figure 1.
-We observe that LLMs are inclined to label two SQL queries are equivalent when they appear similar in structure.
-As a result, it is crucial for LLMs to be able to first spot the differences between these two queries and then apply this knowledge in considering if these two SQL queries are relaxed equivalent or not.
+We observe that LLMs are inclined to label two SQL queries as equivalent when they appear similar in structure.
+As a result, it is crucial for LLMs to be able to first spot the differences between these two queries and then apply this knowledge in considering if these two SQL queries are relaxed equivalent.
 4 Evaluations
-We evaluate the capability of popular large language models in identifying semantic and relaxed equivalence between a pair of SQL queries.
+We evaluate the capabilities of popular large language models in identifying semantic and relaxed equivalence between a pair of SQL queries.
 In particular, we showcase the experimental results using the popular GPT 3.5 turbo (gpt-3.5-turbo) and GPT 4 turbo (gpt-4-1106-preview) models.
@@
 Hyper-parameter: We use zero demonstrations in our prompts.
-In addition, in the API call, we have set temperature to 0 which implies greedy sampling is used to generate output; The max tokens are set to 3000; and the stopping tokens are set to "Answer:" for all experiments.
+In addition, in the API call, we have set temperature to 0 which implies greedy sampling is used to generate the output; The max tokens are set to 3000; and the stopping tokens are set to "Answer:" for all experiments.
 4.1 Data Source
-Our evaluation was conducted on i) Non-equivalent Queries Chu et al.
-(2017) and ii) Spider cross-domain benchmark Yu et al.
+First, we compare our proposed Miniature&Mull prompt with the CounterExample prompt on i) Non-equivalent Queries Chu et al.
+(2017) and ii) Equivalent Queries Zhong .
+These two data sources include pairs of SQL queries that have been expertly labeled as semantically non-equivalent or equivalent.
+  • Nonequivalent Queries Chu et al.
+(2017): Four challenging and one easy questions, including the COUNT BUG, one exam question from CSE344 at the University of Washington, two bugs users found on SQL solver, and a simple hello world example.
+  • Equivalent Queries Zhong et al. (2020a): A hundred pairs of expert labeled equivalent queries.
+Next, we showcase results on the alignment of our proposed LLM-SQL-Solver judgements with human preferences on semantic and relaxed notions using the Spider benchmark Yu et al.
 (2018).
-The Nonequivalent Queries are originally used to evaluate the performance of Cosette a formal language based SQL solver; however, Cosette does not support many SQL operations such as IN, EXCEPT, etc.; and the Cosette project is no longer maintained.
-We use Nonequivalent Queries to investigate if LLMs can successfully notice the semantic differences (often subtle) between two non-equivalent SQL queries.
-  • Nonequivalent Queries Chu et al.
-(2017): It contains 4 challenging and 1 easy questions, including the COUNT BUG, one exam question from CSE344 (data management) at the University of Washington, two bugs users found on SQL solver, and a simple hello world example.
-The Spider benchmark contains 1034 development questions covering 20 databases in different application domains, and each database contains multiple tables.
-The Spider SQL queries are classified into four levels of complexity: easy, medium, hard, and extra.
-This classification is based on various factors, including the quantity of SQL keywords utilized, the incorporation of nested subqueries, and the usage of column selections and aggregations.
-The official evaluation metrics on the Spider benchmark are execution accuracy and exact string match accuracy.
-Since two semantically equivalent SQL queries can be written in many different ways, execution accuracy is often consider as more important than exact string match.
-In all experiments, we use the open sourced execution match evaluation code from GitHub  [^2] https://github.com/THU-BPM/ChatGPT-sql.
-However, as we argued in previous sections, execution match neither captured the semantic equivalence notion nor the relaxed equivalence notion.
-As a result, we showcase the results of using LLMs to determine if the generated SQL queries are the equivalent or not.
-We used three open-sourced results.
-  • Chat Liu et al. (2023): Generate SQL queries using ChatGPT with zero-shot basic prompt.
-  • DIN Pourreza and Rafiei (2023): Generate SQL queries by decomposing the task into fine grained sub-modules and use the GPT-4 model with few-shot prompting
+The Spider benchmark contains 1034 development questions covering 20 databases in different domains.
+The Spider SQL queries are classified into four levels of complexity: easy, medium, hard, and extrahard (extra).
+We evaluate the judgments of LLM-SQL-Solver against execution accuracy (EX) based on human preferences.
+In all experiments, we use the open sourced EX evaluation code from GitHub  [^1] https://github.com/THU-BPM/ChatGPT-sql.
+  • Chat Liu et al. (2023): Generate queries using ChatGPT with the zero-shot basic prompt.
+  • DIN Pourreza and Rafiei (2023): Generate SQL queries by decomposing the task into fine grained sub-modules and use the GPT-4 model with few-shot prompting.
   • DAIL Gao et al. (2023): The current SOTA open sourced technique on the Spider benchmark. It uses the GPT-4 model with few-shot prompting and incorporates self-consistency Wang et al. (2022).
-4.2 Finding Counter Examples
-Judge Prompt Accuracy
-GPT 3.5 turbo CounterExample 4/5
-Miniature&Mull 5/5 $\uparrow$
-GPT 4 turbo CounterExample 3/5
-Miniature&Mull 5/5 $\uparrow$
-GPT 4 CounterExample 4/5
-Miniature&Mull 4/5 -
-Table 1: Results in Nonequivalent Queries
-We test LLMs ability in deciding if two SQL queries are semantically equivalent.
+4.2 Judging Semantic Equivalence
+Judge Prompt Noneq Eq
+GPT 35 turbo CounterExample .8 .45
+Miniature&Mull 1.0 $\uparrow$ .57 $\uparrow$
+GPT 4 turbo CounterExample .6 .7
+Miniature&Mull 1.0 $\uparrow$ .83 $\uparrow$
+Table 1: Results in determining semantic non-equivalent and equivalent queries.
 Recall, two queries can not be semantically equivalent if there exists a counter example (a database instance) such that two queries produce different outputs.
-We used two different prompting methods Counter Example and Miniature & Mull to compare and contrast the accuracy of LLMs.
-An answer is accurate if the answer states the two SQL queries are not equivalent.
-Judge Data Source Equivalence/Prompt Easy Medium Hard Extra Acc
+We used two different prompting methods Counter Example and Miniature & Mull to compare and contrast the accuracy of LLMs in determining semantic equivalence.
+Judge Data Equivalence/Prompt Easy Medium Hard Extra Acc
 - Spider - 248 446 174 166 -
-Execution Match Chat - 221 331 101 72 70.1%
+EX Chat - 221 331 101 72 70.1%
 DIN - 228 382 132 101 81.5%
@@
 DAIL Relaxed/Explain&Compare 220 382 141 110 82.5%
-Table 2: A breakdown of the accuracy scores when using different metrics, i.e., execution accuracy and LLMs judgements with varying prompt techniques.
+Table 2: A breakdown of the accuracy scores when using different metrics, i.e., EX and LLMs’ judgements with varying prompt techniques.
 The compared SQL queries are the gold SQL queries from Spider and the generated SQL queries from Chat, DIN, and DAIL.
-As shown in Table 1, when using the Counter Example prompt (asking LLMs to determine SQL equivalence and provide a counter example if two SQL queries is not equivalent), GPT 4 and GPT 4 turbo failed to notice the subtle difference in the COUNT BUG, whereas GPT 3.5 turbo correctly output that queries in COUNT BUG are not semantically equivalent.
-In addition, GPT 4 turbo failed on a user found bug originally posted on Hackers News  [^3] https://github.com/uwdb/Cosette/issues/29.
+As shown in Table 1, we can observe that the Miniature & Mull prompt always provides a higher accuracy than the Counter Example prompt.
+When using the Counter Example prompt (asking LLMs to determine SQL equivalence and provide a counter example if the SQL queries are not equivalent), GPT 4 turbo failed on a user found bug originally posted on Hackers News lukaseder .
 This bug contains two SQL queries where one used DISTINCT in selection and the other used UNION ALL to union two sub-queries.
-However, GPT 4 turbo failed to recognize that UNION ALL will not remove duplicate rows (though UNION will).
+Hence, GPT 4 turbo failed to recognize that UNION ALL will not remove duplicate rows (though UNION will).
 GPT 3.5 turbo failed on the CSE344 exam question which involves different sub-queries.
-When using Miniature & Mull (asking LLMs to first run a simple example over the given SQL queries and then examining if the example database can be modified to be a counter example), GPT 3.5 turbo and GPT 4 turbo successfully identified all the mistakes.
-However, GPT 4 still failed on the COUNT BUG.
-We can observe that, when using Counter Example prompt (follows the main ideas of chain of thoughts), LLMs scores at least 60% accuracy.
-When using the Miniature & Mull prompting, the accuracy is significantly improved to 100% for GPT 3.5 turbo and GPT 4 turbo.
-As a result, we believe that letting LLMs execute the SQL queries on a simple database instance and then considering a counter example is an effective way to determine SQL semantic equivalence.
-Moreover, since GPT 4 turbo showcases strong performance with faster response time and less monetary costs [^4] https://openai.com/pricing compared to GPT 4, we use GPT 3.5 turbo and GPT 4 turbo for all the following experiments.
+When using Miniature & Mull (asking LLMs to first run a simple example over the given SQL queries and then examining if the example database can be modified to generate a counter example), GPT 3.5 turbo and GPT 4 turbo both successfully identified all the mistakes.
+Our proposed Miniature & Mull prompt not only increased the accuracy of identifying non-equivalent SQL queries to 100% for both models but also increased the accuracy for identifying equivalent SQL queries by at least 12%.
+As a result, we believe that letting LLMs execute SQL queries on a simple database instance and then assessing whether a counterexample can be generated by modifying the database is a more effective method to determine SQL semantic equivalence.
 4.3 Judging NL2SQL Generation
-There are two notions that one can take in judging the equivalence of two SQL queries, i.e., the semantic and relaxed notions.
+There are two notions that one can use in judging the equivalence of two SQL queries, i.e., the semantic and relaxed notions.
 We use the proposed Miniature & Mull and Explain & Compare prompts to evaluate the equivalence of SQL queries following semantic and relaxed notions respectively.
 In Table 2, we present the breakdown of the alignment between human preference and the execution match/LLM judgement on the SQL pairs (gold SQL from Spider and generated SQL from Chat, DIN, and DAIL).
-We observe that the relative ranking between Chat, DIN, and DAIL are preserved for both execution matches and LLM judgements.
+We observe that the relative ranking among Chat, DIN, and DAIL are preserved for both EX and LLM judgements.
 Since the generated SQL queries in Chat are from ChatGPT with zero-shot prompting, we expect the SQL queries from DIN and DAIL to achieve better accuracy.
@@
 As shown in Table 2, the ranking among Chat, DIN, and DAIL based on execution accuracy, GPT 3.5 turbo judgments, and GPT 4 turbo judgements with varying prompts, all aligned with our expectation.
-4.3.1 Agreement between LLMs Judgements and Execution Match
-We observe that for a majority of the questions, LLMs judgements and execution matches agree with each other.
-In particular, as shown in, Table 3, agreements between LLMs and execution match for the semantic equivalence notion have at least 69.6%; similarly, the agreements between LLMs and execution match for the relaxed equivalence notion have at least 82.4%.
-We believe this is an encouraging signal that implies the good quality of execution matches and the good quality of LLMs judgements.
-Assuming the LLMs semantic equivalence judgement is highly accurate, a high percentage of agreements indicate the execution match correctly identified many semantically different SQL queries and the underlying database contains high quality data points (up to 1000 human hours were spent on creating the Spider Benchmark Yu et al.
+4.3.1 Agreement between LLMs’ Judgements and Execution Accuracy
+For a majority of the questions, LLMs’ judgements and EX agree with each other.
+In particular, as shown in, Table 3, agreements between LLMs and EX for the semantic equivalence notion have at least 69.6%; similarly, the agreements between LLMs and EX for the relaxed equivalence notion have at least 82.4%.
+We believe this is an encouraging signal that implies the good quality of the Spider benchmark and LLMs’ judgements.
+The Spider benchmark contains high quality data points (1000 human hours spent Yu et al.
 (2018)).
-If the underlying database does not contain enough high quality data points, then many of the non-semantically equivalent SQL queries will return the same results.
-Assuming the underlying database is well designed and the execution match is highly accurate, then a high agreement between execution match and LLMs are obviously desirable.
+Creating these high quality databases, which is necessary for good EX, are expensive and does not scale.
+Assuming the underlying database is well designed and the EX is highly accurate, then a high agreement between EX and LLMs are obviously desirable.
 Judge Equiv Data Agreement
@@
 DAIL 86.1%
-Table 3: We compare the agreement between LLMs judgements and execution matches.
+Table 3: We compare the agreement between LLMs’ judgements and execution matches.
 4.3.2 Human Evaluation on Samples
-Although, the agreements between execution accuracy and LLM’s judgments are high, there is a still lots of questions with disagreements.
-In this section, we provide a detailed analysis on the sampled questions from the set of questions where execution match and LLM’s judgement differs.
-Since there are hundreds of such questions (different results between execution match and LLM) per data source and per GPT models, we sample 10 questions randomly from the SQL query pairs where LLMs and execution match disagree.
+Although, the agreements between execution accuracy and LLM’s judgments are high, there are still lots of questions with disagreements.
+In this section, we provide a detailed analysis on the sampled questions from the set of questions where EX and LLM’s judgement differs.
+Since there are hundreds of such questions (different results between EX and LLM) per data source and per GPT models, we sample 10 questions randomly from the SQL query pairs where LLMs and EX disagree.
 Equivalence Evaluated Manually To evaluate the equivalence between two SQL queries, we use the human preference approach.
-Three of the authors manually check the equivalence of the sampled questions where LLMs and execution match disagree.
+Three of the authors manually checked the equivalence of the sampled questions where LLMs and EX disagree.
 In addition, we have also differentiated the equivalence between the semantic notion and relaxed notion (in essence, the former is objective and the latter is subjective).
-To ensure the quality of human judgement, the annotations are conducted independently, and we use the majority vote to determine the label for each pair of SQL queries (i.e., each pair of SQL queries are examined by three times independently).
+To ensure the quality of human judgement, the annotations are conducted independently, and we use the majority vote to determine the label for each pair of SQL queries.
 Judge Equiv Data Prefer
@@
 DAIL 7/10 $\uparrow$
-Table 4: We compare the human preference between LLM judge and execution match on 10 randomly sampled questions from different data sources.
-The questions are selected from a set where LLMs and execution match have different opinions on these questions
+Table 4: We compare the human preference between LLM judge and EX on 10 randomly sampled questions from different data sources.
 Semantic Equivalence
-When using the semantic equivalence notion, we find GPT 3.5 turbo may not deliver better judgment than execution match, but GPT 4 turbo showcase better alignments.
-When the execution match detect discrepancy in the SQL queries’ outputs, it often indicates these two SQL queries are not semantically equivalent.
-There are a few rare cases that we observed which the SQL queries are semantically equivalent but execution match determines not equivalent.
-For example, when the provided database in Spider contains column values that can not be decoded (similar issues are also raised  [^5] https://github.com/taoyds/spider/issues/46), semantically equivalent queries can be misclassified by execution match.
-Figure 2: The correct operation rates for LLMs. Correct operations imply the LLM correctly executed and reasoned about the SQL queries on its own provided simple database.
+When using the semantic equivalence notion, we find GPT 3.5 turbo may not deliver better judgment than EX, but GPT 4 turbo results in better alignments.
+When EX detects discrepancy in the SQL queries’ outputs, it often indicates these two SQL queries are not semantically equivalent.
+There are a few rare cases that we observed in which the SQL queries are semantically equivalent but EX determines not equivalent.
+For example, when the provided database in Spider contains column values that can not be decoded (similar issues are also raised sythello , semantically equivalent queries can be misclassified by EX.
+Figure 2: The rates of LLM correctly executed the SQL queries on its own provided simple database.
 In Table 4, GPT 3.5 turbo aligned with human experts in the range of two to five out of ten.
-Diving deep into the questions which GPT 3.5 turbo wrongly classified the SQL queries, we find that there are common reasons for the wrong classifications: Distinct entries vs duplicated entries.
+Diving deeper into the questions which GPT 3.5 turbo wrongly classified the SQL queries, we find that there are common reasons for the wrong classifications: Distinct entries vs duplicated entries.
 GPT 3.5 turbo often failed to consider the difference between selecting unique rows and selecting rows that may be duplicated; Subtle difference in column names.
-When the gold SQLs and generated SQLs share very similar SQL structure but contains subtle differences in the columns used such as group by id v.s.
-group by student_id, GPT 3.5 turbo may fail to notice the small difference; Hallucination.
-GPT 3.5 turbo may returns wrong results from executing the SQL queries on its own created simple database instance and sometimes when the outputs are indeed identical, it may still state that the outputs are different.
+When the gold SQLs and generated SQLs contain only subtle differences in the columns used such as group by id v.s.
+group by student_id, GPT 3.5 turbo may fail to notice these differences; Hallucination.
+GPT 3.5 turbo may return wrong results when executing the SQL queries on its own created simple database instance and sometimes when the outputs are indeed identical, it may still state that the outputs are different.
 This hallucination snowballing effect Zhang et al.
-(2023a) is one of the top reason why GPT 3.5 turbo perform worse compared to GPT 4 turbo.
-When GPT 4 turbo is used as the judge, GPT 4 turbo constantly produced more accurate or comparable results to execution match, with at least 5 out of 10 questions for DAIL and at most 10 out of 10 questions for Chat.
+(2023a) is one of the top reason why GPT 3.5 turbo performs worse compared to GPT 4 turbo.
+When GPT 4 turbo is used as the judge, it constantly produced more accurate or comparable results to EX, with at least 5 out of 10 questions for DAIL and at most 10 out of 10 questions for Chat.
 One interesting observation is that GPT 4 turbo correctly identified the difference between order by asc/desc limit 1 v.s.
@@
 The latter, however, may return many rows when these rows share the same min or max value.
-Execution match often failed to capture this difference.
-Although GPT 4 turbo judgment is much better than GPT 3.5 turbo, GPT 4 turbo may still miss on the small differences such as similar but different column names (such as song_name v.s.
-name), overlook the fact that extra columns are selected, and also sometimes hallucinate.
-As shown in Figure 2, we looked into how well does LLMs executed and reasoned about the SQL queries on its own generated simple database.
-We find that GPT 4 turbo has much higher correct operation rates which leads much better response comapred with the weaker model.
-One common phenomena that lead to false answers in execution match is due to the return of the empty tuple.
+EX often failed to capture this difference.
+Although GPT 4 turbo’s judgment is much better, GPT 4 turbo may still miss some minor differences such as similar but different column names (such as song_name v.s.
+name), overlook the fact that extra columns are selected, and also sometimes hallucinates.
+As shown in Figure 2, we explored how well do LLMs execute and reason about the SQL queries on their own generated simple database.
+GPT 4 turbo has much higher correct operation rates which leads to much better responses compared with the weaker model.
+One common phenomena that lead to false answers in EX is due to the return of the empty tuple.
 When the underlying database does not contain tuples that match the predicate, gold SQLs and generated SQLs will always match, irrespective of how different they are.
 Another interesting observation raised from a few comparisons is about the case sensitivity.
-OpenAI GPT models think databases are by default case insensitive; Spider benchmark, on the other hand, uses the SQLite databases which are case sensitive by default.
-Many popular database are case insensitive such as SQL Server Mistry and Misner (2014) and MySQL DuBois (2013).
-We believe that treating case sensitivity in string comparison as an excessive step, and therefore, we operate under the assumption that the underlying database instance is case insensitive when evaluating semantic equivalence.
-Relaxed Equivalence When using the explain and compare prompt for determining the relaxed equivalence, both GPT 3.5 turbo and GPT 4 turbo are consistently better aligned with human judgment or at par compared to execution match, as shown in Table 4.
-Recall, the relaxed notion aims to illustrate the degree to which humans perceive two queries as being similar in their meaning, and mimics the opinion of the users of NL2SQL services.
-The positive results on the alignment between LLMs judgments with human preferences indicate optimizing for LLMs judgments accuracy is a better optimization target than execution match accuracy.
-The main reasons for why LLMs judgements are better aligned with human preference than execution match is that LLMs can zoom into the query logic and determine whether the two SQL queries share the same logic and contain the same truth.
+OpenAI GPT models consider databases are by default case insensitive; Spider benchmark, on the other hand, uses the SQLite databases which are case sensitive by default.
+However, many popular database are by default case insensitive such as SQL Server Mistry and Misner (2014) and MySQL DuBois (2013).
+Therefore, we operate under the assumption that the underlying database instance is case-insensitive when evaluating semantic equivalence.
+Relaxed Equivalence.
+When using the Explain & Compare prompt for determining the relaxed equivalence, LLMs are consistently better aligned with human judgments than EX, as shown in Table 4.
+The relaxed notion aims to illustrate the degree to which humans perceive two queries as being similar in their meaning, and mimics the opinions of users.
+The positive results on the alignment between LLMs’ judgments with human preferences indicate optimizing accuracy for LLMs’ judgments is a better optimization target than EX.
+The main reasons for why LLMs’ judgements are better aligned with human preference than EX is that LLMs can zoom into the query logic and determine whether the two SQL queries share the same logic and contain the same truth.
 Consider the following example, order by col asc/desc limit 1 v.s.
 where col = select min/max(col).
-They may return different or the same outputs after execution (depends on the underlying database instance).
-Since execution match strictly checks the outputs, it may sometimes return equivalent and other times return not equivalent.
-However, LLMs can correctly identify that these two queries are selecting the rows with minimal or maximal value and these two queries are equivalent under the relaxed notion,
+LLMs can correctly identify that these two queries are selecting the rows with minimal or maximal value.
+Since EX strictly checks the outputs, it may sometimes return equivalent and other times return not equivalent.
 One interesting observation we found is about null values.
 GPT often consider COUNT(*), COUNT(primary key), and COUNT(DISTINCT primary key) to be different.
-The reason is that GPT models think that null values can be inserted into the column with primary key constraint.
-In fact, the SQLite databases [^6] https://www.sqlite.org/lang_createtable.html, by default, allow null values to be inserted to primarily key.
-However, in many database primary key constraints explicitly states that null values are not allowed [^7] https://docs.oracle.com/javadb/10.10.1.2/ref/rrefsqlj13590.html.
+The reason is that GPT models consider that null values can be inserted into a column with primary key constraint.
+In fact, the SQLite databases, by default, allow null values to be inserted to primary key column.
+However, in many database primary key constraints explicitly states that null values are not allowed Oracle .
 Based on the relaxed equivalence definition, we consider records in the primary columns to be not null.
-Although LLMs correctly argued that when null values are presented in the primary key column, then the given SQL queries lead to different results (such as COUNT(*) considers null values as unique entries, where as COUNT(primary key) ignores rows with null values), we consider these not as a significant issue.
+Although LLMs correctly consider that having null values in the primary key column may lead to different results (note that COUNT(*) considers null values as unique entries, where as COUNT(primary key) ignores rows with null values), we consider these not as a significant issue.
 5 Discussion and Limitation
+In this discussion section, we would like to highlight the broader implications and potential integration of our proposed LLM-SQL-Solver.
+During the SQL rewrite phase of the query optimization engine, the input SQL query is transformed into a new SQL query that can be optimized more easily.
+The key challenge in query rewriting is to determine if the input query and the rewritten query are indeed semantic equivalent.
+The LLM-SQL-Solver can be directly used as a tool, during the rewriting phase, to check semantic equivalence and ensure bugs (e.g., COUNT bug Ganski and Wong (1987)) do not occur in the data system.
+Furthermore, the integration of the LLM-SQL-Solver as a text-to-SQL evaluation metrics using the relaxed equivalence notion could improve the overall quality of user experience and enable more human-like capabilities.
+Concurrent with our research, the chain-of-table approach Wang et al.
+(2024) independently explores a similar concept, allowing LLMs to form a reasoning path through table modifications to improve question answering from tabular data, akin to our Miniature & Mull prompt.
+However, while our technique focuses on determining query semantic equivalence through table creation and modification, their goal is to deepen the understanding of the tabular data itself.
 There are some limitations or areas of improvement to this work.
-First, we only evaluated the capability of Large Language Models of OpenAI.
-There are more other models such as Llama2 7B Touvron et al.
+First, we only evaluated the capabilities of the Large Language Models of OpenAI.
+There are other alternative models such as Llama2 7B Touvron et al.
 (2023), Vicuna Chiang et al.
@@
 (2023b).
-Nevertheless, we believe that GPT 3.5 turbo and GPT 4 turbo exemplify the current status of GPT model family’s capabilities.
-In addition, in the manually evaluating the equivalence under two different views (semantic and relaxed), it is done over the sampled questions.
-Since there are more than two thousand of SQL queries per data source and equivalence view, manually checking all of them are arguably infeasible.
+Nevertheless, we believe that GPT 3.5 turbo and GPT 4 turbo exemplify the current status of the GPT model family’s capabilities.
+In addition, our evaluations on Spider Benchmark were conducted over randomly sampled questions, in which we have to manually label the equivalence of the two different notions (semantic and relaxed).
+Since there are nearly ten thousand pairs of SQL queries, manually checking all of them are arguably infeasible.
 In fact, even checking a pair of SQL queries for semantic view takes around 10 minutes for experts in SQL.
@@
 6 Conclusion
-In this paper, we studies and answer two main questions: i) Can current LLMs be used to help data engineers in checking semantic equivalences between SQL queries and ii) should we use LLMs to judge the quality of generated SQL queries for text-to-SQL benchmarks.
+In this paper, we studied and answered two main questions: i) Can current LLMs be used to help engineers in checking semantic equivalences between SQL queries and ii) should we use LLMs to judge the quality of generated SQL queries for text-to-SQL benchmarks.
+To the best of our knowledge, this is the first work studying the ability of LLMs in understanding SQL queries and determining SQL equivalences.
 Through extensive evaluations and manually examining the quality of the answers based on the alignment with human preferences, we find current LLMs (especially GPT 4 turbo), using our proposed prompt techniques, can assists engineers in determining semantic equivalent queries, but may yield incorrect responses.
-On the other hand, when comparing the generated SQL query with the gold SQL query with relaxed notion of equivalence, LLMs’ responses constantly achieve high alignments with human preferences.
-Moreover, another advantage of using LLMs to determine SQL equivalence is that it can avoid the construction of sophisticated databases, and no data stored need to be accessed.
-To the best of our knowledge, this is the first work studying the ability of LLMs in determining SQL equivalence.
-As the learning ability in LLMs continuously to advance, we believe that using these models to check SQL equivalence is a promising direction.
+On the other hand, when comparing the generated SQL query with the gold SQL query with our relaxed notion of equivalence, LLMs’ responses constantly achieve high alignments with human preferences.
+Moreover, our proposed LLM-SQL-Solver avoids accessing stored user data, and hence no sophisticated databases need to be constructed.